A 3.10 m uniform ladder weighing 8- N is placed against a smooth wall so that its lower end is 2 m away from the wall. (a) What is the force exerted by the ladder against the wall? (b) What is the direction and magnitude of the force exerted by the ladder against the floor?


1
Expert's answer
2021-07-14T10:14:58-0400
"Fl\sin{\alpha}=0.5lW\cos{\alpha}\\\alpha=\arccos{\frac{2}{3.1}}=49.8\degree\\F\sin{49.8}=0.5(8)\cos{49.8}\\F=3.38\ N"